a loan of $500 is to be paid with two payments of $300, one in 3 months, and another in 6 months. the compound interest charged per annum on the loan is;
1
Expert's answer
2020-06-25T10:57:18-0400

"0=(500*(1+\frac{i}{12})^3-300)*(1+\frac{i}{12})^3-300"

"(1+\frac{i}{12})^3=t"

"0=500*t^2-300*t-300"

"t=\frac{3-8.31}{10}" - less than 0, interest rate can't be less than 0

"t=\frac{3+8.31}{10}=1.131"

"(1+\frac{i}{12})^3=1.131"

"i=0.5027=50.27\%"

i-annual interest rate